OpenStudy (anonymous):

radius and interval of convergence of the series sqrt(n)*x^n as n goes from 1 to infinity?? please help..

OpenStudy (anonymous):

the radius = 0, since x stands alone here :)

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

Just give me a minute.

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

are you sure it's the sqrt of n and not the power of n?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well.. I am going to use the ratio test first to find the radius of convergence.

OpenStudy (anonymous):

you can't use the ratio test here anwar, since it's to the power of 1/n and not n, you'll have to use the root test here ^_^ and hey :)

OpenStudy (anonymous):

but since the question doesn't have any sort of exponential function, then "maybe" you can :)

OpenStudy (anonymous):

but it's going to be difficult

OpenStudy (anonymous):

Yes I can, and it won't be difficult. And hello! :)

OpenStudy (anonymous):

lol, root test is much easier here :)

OpenStudy (anonymous):

let: \[an = x^n\] so \[|an^{\frac{1}{n}} |= (|x^n)^{\frac{1}{n}} |= |x|\]

OpenStudy (anonymous):

and :\[0\le |x| < 1\] so the radius is = 1 in this case, I guess, correct me if I'm wrong please :)

OpenStudy (anonymous):

so since :\[0 \le |x| < 1\] then : \[-1 < x < 1\] is the interval of convergence ^_^

OpenStudy (anonymous):

I took this section yesterday morning during my calculus class lol

OpenStudy (anonymous):

so I might have a couple of mistakes ^_^" maybe

OpenStudy (anonymous):

just so its clear the series is..\[\sqrt{n}*x ^{n}\]

OpenStudy (anonymous):

oh >_<! I thought it was to the power of 1/n

OpenStudy (anonymous):

\[\left| {a_{n+1} \over a_n} \right|=\left| {\sqrt{n+1}. x^{n+1} \over \sqrt{n} x^n} \right|=\left| x \right|.\left| {\sqrt{n+1} \over \sqrt n} \right|\]

OpenStudy (anonymous):

then find the limit , oh my bad, I completely misunderstood the question >_<

OpenStudy (anonymous):

proceed with anwar's answer, it seems correct :)

OpenStudy (anonymous):

oh sorry for the confusion..ok yea i got till there..and then?

OpenStudy (anonymous):

continue anwar, you're on the right track :)

OpenStudy (anonymous):

So now take the limit as n goes to infinity: \[\lim_{n \rightarrow \infty}{\left| x \right|.\left| {\sqrt{n+1} \over \sqrt n} \right|}=\left| x \right|\]

OpenStudy (anonymous):

oh, almost the same LOL =P bleh

OpenStudy (anonymous):

Is the limit part clear?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

oh congrats on being a champion anwar ^_^

OpenStudy (anonymous):

Good. As I said before, the series is convergent when the limit value is <1. In our case, the radius of convergent is: \[R=\left| x \right|<1 \implies -1<x<1\]

OpenStudy (anonymous):

Thanks sstarica :)

OpenStudy (anonymous):

:) sure

OpenStudy (anonymous):

the radius of convergence = 1 and the interval is the one on top lol ^_^

OpenStudy (anonymous):

Now you should check the endpoints. check if the series converges at x=1 and x=-1.

OpenStudy (anonymous):

but i thought it had to be less than one to converge..

OpenStudy (anonymous):

oo nevermind. understood

OpenStudy (anonymous):

Well.. It converges "for sure" when it's less than 1, and diverges when it's greater than one. But the test fails when it's equal to 1. So we have to check it separately. Does that make sense?

OpenStudy (anonymous):

does x also have to be greater than -1?? y is that?

OpenStudy (anonymous):

Yeah. because we had IxI<1, that means x is between -1 and 1.

OpenStudy (anonymous):

can we always use the ratio test for this and if we do that isnt x always in between -1 and 1?

OpenStudy (anonymous):

I have an exam today, and still have a lot of material to study. So, I should get going now :(. I am sure sstarica will be here to help.

OpenStudy (anonymous):

oh ok then! thanks for all the help!

OpenStudy (anonymous):

No, you don't have always to use the ratio test. You can use other tests as well, especially the root test. It just depends on the series itself. The radius of convergence differs from series to another. GOOD LUCK!

OpenStudy (anonymous):

http://www.youtube.com/watch?v=01LzAU__J-0

OpenStudy (anonymous):

thank u so much! good luck to you too!